Use differentials to approximate the volume of material needed to make a rubber ball if the radius of the hollow inner core is 2 in., and the thickness of the rubber is ⅛ in.

1
Expert's answer
2021-11-23T17:59:53-0500

volume of the sphere is "V=\frac{4}{3}\pi r^3"

"dV=\frac{4}{3}\pi (3r^2)dr"

"=4 \pi r^2dr\\=4\pi (4)^2(0.2)=12.8 \pi"